Speaker’s Name

About Our Next Speaker

Current affiliation of Speaker’s Name

Professor of mathematics, University of Wherever. Junior partner at company X. Speaker for organization/project X.

Experience and achievements

Academic degree, but only if appropriate Current and/or previous positions, possibly with dates Publications (possibly just number of publications) Awards, prizes

Concerning today’s talk

Expert who has worked in the field/project for X month/years. Will present his/her/group’s/company’s research on the subject. Will summarize project report or current project status.